Four equal-magnitude point charges (3.0 C) are placed at the corners of a square that is 40 cm on a side. Two, diagonally opposite each other, are positive, and the other two are negative. Find the force on either negative charge.

1
Expert's answer
2022-01-26T07:39:59-0500

"F=2F_1+F_2=2\frac{kq^2}{r^2}-\frac{kq^2}{2r^2}=\frac{3kq^2}{2r^2}=760\cdot 10^9~N."